OpenStudy (anonymous):

can you explain me newton's 2nd law?

OpenStudy (anonymous):

the acceleration produced in an object is directly proportional to the force applied

OpenStudy (anonymous):

@him is correct. 1st:objects in motion tend to stay in motion and vece verse, 2nd: forces come in equal and opposite pairs, 3rd: Forces produce accelerations directly proportional to the mass of the object being accelerated or F = ma. Actually, forced produce changes i momentum, but its essentially the same thing. If you have any questions just ask :)

OpenStudy (anonymous):

Newton's second law of motion pertains to the behavior of objects for which all existing forces are not balanced. The second law states that the acceleration of an object is dependent upon two variables - the net force acting upon the object and the mass of the object. The acceleration of an object depends directly upon the net force acting upon the object, and inversely upon the mass of the object. As the force acting upon an object is increased, the acceleration of the object is increased. As the mass of an object is increased, the acceleration of the object is decreased. Newton's second law of motion can be formally stated as follows: The acceleration of an object as produced by a net force is directly proportional to the magnitude of the net force, in the same direction as the net force, and inversely proportional to the mass of the object. a = F(net)/m

OpenStudy (anonymous):

If u visit the following site, it will make things very clear for you.... http://zonalandeducation.com/mstm/physics/mechanics/forces/newton/newtonLaw2.html

OpenStudy (anonymous):

an expression for how an object will accelerate based on its mass and the net force applied to the object. This law can be summarized by the equaon: ΣF = m·a

OpenStudy (anonymous):

the force applied on a body(at motion) is equal to the rate of change of momentum if p(momentum)=mv(mass*velocity) f=dp/dt f=m*(dv/dt)+v*(dm/dt)(second part becomes 0 ,change in mass is at high velocities only ) so F=ma(where 'a' is the acceleration)